часот отр For 2019, SoakNFun expects a sales mix of two individual tickets for every three family tickets. SoakNFuns total fRequirement 1. Compute the weighted average contribution margin per ticket. Complete the table below to calculate the weighteRequirement 1. Compute the weighted-average contribution margin per ticket. Complete the table below to calculate the weighteRequirement 2. Calculate the total number of tickets SoakNFun must sell to break even. Start by selecting the formula and entVVLIVUILTUUveluyCULLI IULIULUI Y P ULL Requirement 2. Calculate the total number of tickets SoakNFun must sell to break even.Requirement 3. Calculate the number of individual tickets and the number of family tickets the company must sell to break eve
Answer #1

1.

Individual Family Total
Sale price per ticket $                  35 $                  105
- Variable cost per ticket $                  25 $                  100
Contribution Margin per unit $                  10 $                      5
x Sales mix in units 2 3
Contribution Margin $                  20 $                    15 $                     35
Weighted-average Contribution Margin per unit $                       7

2. Required sales in units = (Fixed Costs + Target Income) / Weighted-average contribution margin per unit
= ($19250+0)/7 = 2750 units

3. Break Even Sales = Number of tickets to break even x Proportion
Individual = 2750 x 2/5 = 1100 tickets
Family = 2750 x 3/5 = 1650 tickets